OpenStudy (anonymous):

An automobile moving along a straight track changes its velocity from 40 m/s to 80 m/s in a distance of 200 m. What is the (constant) acceleration of the vehicle during this time? Question 13 options: 1) 8.0 m/s 2) 9.6 m/s 3) 12 m/s 4) 6.9 m/s 5) 0.20 m/s

OpenStudy (turingtest):

Fundamental kinematic equation:\[v_f^2=v_i^2+2ad\to a=\frac{v_f^2-v_i^2}{2d}\]now just plug in the numbers

OpenStudy (anonymous):

3 is correct from independent equation motion we have it